# Angular Momentum Drain: Despinning Embedded Planetesimals

### v1

@@ 4.1 A Simple Velocity Gradient
Using Equations 43 & 46, we solve for the magnitude of the linear velocity gradient, $\gamma$:
@@ 4.2 Effective Spindown Over the Lifetime of the Collisional Era
We adopt the circumstellar disk model from Quillen et al. (2024) which has the power law form for temperature and surface density of the minimum mass solar nebula by Hayashi (1981)
$\Sigma_{g}$ $=g_{\Sigma}\left(\frac{r}{10\text{AU}}\right)^{-3/2}$ (51)
$T_{g}$ $=g_{T}\left(\frac{r}{1\text{AU}}\right)^{-1/2}$ (52)
where $g_{\Sigma}$ and $g_{T}$ are coefficients that can be used to adjust disk surface density and temperature. For this disk model the headwind velocity is independent of radius and $u_{hw}\approx 38$ m/s.
@@
$\approx 8\times 10^{-9}\ {\rm kg~{}m}^{-2}{\rm s}^{-1}\left(\frac{f_{p}}{10^{-2}}\right)$
×$\ \ \ \times\left(\frac{r}{10\ {\rm AU}}\right)^{-\frac{11}{4}}\frac{b_{max}^{2}}{R_{a}^{2}}g_{\Sigma}g_{T}^{\frac{1}{2}}.$ (53)
Assuming a planetesimal is embedded in disk undergoing streaming instability for $10^{3}$ years at 10 AU, with $f_{p}\sim 10^{2}$ (as streaming instability concentrates the particles), we estimate a total mass in projectiles
@@
$\sim 3\ \left(\frac{f_{p}}{10^{2}}\right)\left(\frac{r}{10\ {\rm AU}}\right)^{-\frac{11}{4}}g_{\Sigma}g_{T}^{\frac{1}{2}}\left(\frac{\rho_{a}}{235\ {\rm kg~{}m}^{-3}}\right)^{-1}$
×$\ \ \ \ \times\left(\frac{\Delta t}{10^{3}\ {\rm yr}}\right)\left(\frac{R_{a}}{10\rm km}\right)^{-1}\left(1+\frac{2GM_{a}}{u_{hw}^{2}R_{a}}\right).$ (55)
During an epoch of streaming instability, a recently formed planetesimal could be impacted by about its own weight in pebbles. Nevertheless because escaping ejecta is inefficient at removing spin angular momentum, at most a few percent of its spin could be carried away by ejecta even with this high impact rate.
In this section we found that the total mass in impacts experienced by an embedded planetesimal could be large compared to its total, particularly if the planetesimal is small. In our study we have found that impacts can drain a planetesimal’s angular momentum, however an impact more conventionally deflects an object. Computation of the momentum transfer efficiency would require taking into account the different directions that ejecta escape the planetesimal, during gravitational defocusing, as discussed in A. As the ejecta velocity is slower than the impact velocity, gravitational defocusing could be an even stronger effect on the escaping ejecta than on the incoming projectile.
@@ 4.3 Caveat: Impactor Density Distribution
Throughout our integrations, we assumed all collisions occurred with a single projectile density $\rho_{pj}$. Since the ratio of projectile to planetesimal density, $\pi_{4}$, affects the integrated erosion and angular momentum transfer efficiency (See Section 4), the impactor density distribution would affect the total spin-down observed, due to the total projectile mass not all being of the same density. As such, evolutionary models on the size and density distribution of disk particulates (e.g. Colmenares et al. 2024) could be used to expand our extend our models.
@@ 5 Discussion
We discuss our work in the context of the study by Nesvorný et al. (2021), who studied KBO binary formation with N-body soft sphere (PKDGRAV). Nesvorný et al. (2021) found that the angular momentum of a pebble cloud prevents the rapid collapse into a single planetesimal. Instead, binary systems were more likely to form, with only a fraction of the pebble cloud being incorporated into the planetesimals. In this section we consider the fraction of pebble cloud angular momentum that must be removed to collapse a single planetesimal and place this value in context with the fraction of angular momentum that can be removed by escaping ejecta.
@@ 5.1 The Angular Momentum of Planetesimals Formed Via Streaming Instability
Equation 59 implies that the problem of angular momentum removal is worse in the outer solar system than in the inner solar system. In the outer solar system, due to the large fraction of angular momentum that must be lost, the angular momentum drain mechanism investigated in this paper would be incapable of facilitating collapse into a single planetesimal.
@@ 6 Summary
We find that the angular momentum drain mechanism proposed by Dobrovolskis and Burns (1984) operates on planetesimals that are embedded within a circumstellar disk and are experiencing impacts originating from a headwind. We characterize the efficiency of the process by computing the ratio of the angular momentum lost per unit projectile mass and the planetesimal’s spin angular momentum. We find that the angular momentum drain mechanism is inefficient, with efficiency $\epsilon_{L}\ll 1$ and at most a few percent. Thus, if the impacting material is associated with a pebble cloud and contains angular momentum, the impact mediated angular momentum drain mechanism, would not significantly facilitate planetesimal collapse. With N-body simulations Nesvorný et al. (2021) found that the angular momentum of a pebble cloud prevents the collapse into a single planetesimal. Our study supports this result but with a different numerical approach. In our study the angular momentum transfer arises through ejecta scaling rather than dissipation associated with inelastic collisions.
We find the angular momentum transfer efficiency is most sensitive to headwind velocity and the ratio of planetesimal to projectile density. As planetesimal mass increases, the effects of gravitational focusing cause a decrease in the angular momentum transfer efficiency and an increase in the mass loss ratio due to the higher impact velocity caused by gravitational focusing. Assuming the projectiles originate from a disk filament, that might be present due to streaming instability with a velocity gradient capable of generating the computed pebble cloud, we find no meaningful reduction of the angular momentum transfer efficiency with regard to spinning down an embedded planetesimal. Future studies could take into account the angular momentum that is inherent in a population of projectiles.
The impact related mechanism of angular momentum loss, because it is inefficient, does not help resolve the issue of how excess angular momentum is shed in a collapsing pebble cloud. Additional sources of angular momentum drain may be necessary for the collapse of single-planetesimals. Pebble cloud collapse could lead preferentially to binary formation, as explored by Nesvorný et al. (2019, 2021). Alternatively, planetesimal formation itself could be inefficient with many pebble clouds failing to collapse, and only those that do collapse initially having low angular momentum.
@@
The python code used in this study is archived at http://xxxxarxiv.org will be updated upon posting to arxiv
@@ Appendix A The Momentum and Angular Momentum Transfer
Assuming that the planetesimal is experiencing many collisions, if the distribution of projectiles has no angular momentum then the $m_{pj}{\bf r}\times{\bf u}_{pj}$ term in equation 67 will average to zero. We note that in equations 64 and 67, the relevant escaping ejecta momentum is computed in the frame rotating with the surface at the site of impact.
Up to this point, we have neglected gravitational focusing. Equations 60 and 61 describe the impact and ejecta curtain just prior to and after the impact. The projectile velocity ${\bf u}_{pj}$ refers to velocity of the projectile relative to the planetesimal at the moment of impact (in an inertial, not rotating frame) and is not equal to the velocity of the projectile at large distance from the planetesimal. As a consequence equation 66 describes the momentum transfer efficiency $\beta$ only for high velocity projectiles.

### v2

@@ 5.1 A Simple Velocity Gradient
Using Equations 45 & 48, we solve for the magnitude of the linear velocity gradient, $\gamma$:
@@ 5.2 Effective Spindown Over the Lifetime of the Collisional Era
We adopt the circumstellar disk model from Quillen et al. (2024) which has the power law form for temperature and surface density and is based on a modified version of the minimum mass solar nebula by Hayashi (1981)
$\Sigma_{g}$ $=3430\ {\rm kg\ m}^{-2}\ g_{\Sigma}\left(\frac{a}{10\text{AU}}\right)^{-3/2}$ (53)
$T_{g}$ $=170\ {\rm K}\ g_{T}\left(\frac{a}{1\text{AU}}\right)^{-1/2}$ (54)
where $g_{\Sigma}$ and $g_{T}$ are ratios that can be used to adjust disk surface density and temperature. Following the disk model from Quillen et al. (2024), these ratios are equal to $1$. For this disk model the headwind velocity is independent of radius and $u_{hw}\approx 38$ m/s.
@@
$\approx 1.28\times 10^{-10}\ {\rm kg~{}m}^{-2}{\rm s}^{-1}\left(\frac{f_{p}}{10^{-2}}\right)$
×$\ \ \ \times\left(\frac{a}{45\ {\rm AU}}\right)^{-\frac{11}{4}}\frac{b_{max}^{2}}{R_{a}^{2}}g_{\Sigma}g_{T}^{\frac{1}{2}}.$ (55)
Assuming a planetesimal is embedded in disk undergoing streaming instability for $10^{4}$ years at 45 AU, with $f_{p}\sim 10^{2}$ (as streaming instability concentrates the particles), we estimate a total mass in projectiles
@@
$\sim 0.52\ \left(\frac{f_{p}}{10^{2}}\right)\left(\frac{a}{45\ {\rm AU}}\right)^{-\frac{11}{4}}g_{\Sigma}g_{T}^{\frac{1}{2}}\left(\frac{\rho_{a}}{235\ {\rm kg~{}m}^{-3}}\right)^{-1}$
×$\ \ \ \ \times\left(\frac{\Delta t}{10^{4}\ {\rm yr}}\right)\left(\frac{R_{a}}{10\rm km}\right)^{-1}\left(1+\frac{2GM_{a}}{u_{hw}^{2}R_{a}}\right).$ (57)
During an epoch of streaming instability, a recently formed planetesimal at 45 AU could be impacted by about half its own weight in pebbles. Depending on the specific projectile and material parameters, this mechanism could result in the removal of 30% - 50% of a planetesimals spin angular momentum. Angular momentum drain can occur in both accretionary regimes as well as erosional regimes, with planetesimal radii greater than 10 km always being accretionary except for $u_{hw}>40$ m/s, and for a strengthless planetesimal, where the accretionary regime begins at slightly larger planetesimal radii. We find that for a strengthless planetesimal, while small-radii is dominated by the effects of oblique impacts (Discussed in greater detail in Section 5.4), at large radii results in an additional $\sim$ 40% angular momentum drain in comparison to planetesimals with some small but non-zero cohesive strength.
In this section we found that the total mass in impacts experienced by an embedded planetesimal is a fraction of its total mass. We additionally found that assuming a velocity gradient across the headwind particles does not influence the angular momentum drain nor the mass loss ratio to a meaningful degree. In our study we have found that impacts can drain a planetesimal’s angular momentum, however an impact more conventionally deflects an object. Computation of the momentum transfer efficiency would require taking into account the different directions that ejecta escape the planetesimal, during gravitational defocusing, as discussed in A. As the ejecta velocity is slower than the impact velocity, gravitational defocusing could be an even stronger effect on the escaping ejecta than on the incoming projectile.
@@ 5.3 Caveat: Impactor Density Distribution
Throughout our integrations, we assumed all collisions occurred with a single projectile density $\rho_{pj}$. Since the ratio of projectile to planetesimal density, $\pi_{4}$, affects the integrated erosion and angular momentum transfer efficiency (See Section 5), the impactor density distribution would affect the total spin-down observed, due to the total projectile mass not all being of the same density. As such, evolutionary models on the size and density distribution of disk particulates (e.g. Colmenares et al. 2024) could be used to expand our models.
5.4 Caveat: Highly Oblique Impacts
Integrating collisions across the planetesimal surface, we expect many impacts to occur at highly oblique impact angles. The crater scaling laws we use are well tested at normal impacts (Housen and Holsapple, 2011), and have been adjusted to account for oblique impacts down to 30 degrees (Raducan et al., 2022). Many of the effects we see occur at these impacts with extreme obliquity. Namely, the effective spin-up for low strength planetesimals, or for large values of $\pi_{4}$, occur due to the ejecta velocity following a butterfly pattern at large obliquities (Impact angles below 20 - 40 degrees). As the planetesimal radii further decrease, more and more material escapes and the variations in the butterfly pattern on each hemisphere due to the effects of rotation fade away as all of the crater mass escapes. In the case for strengthless materials, this effect has not yet reversed at the lowest radii tested ($R_{a}=1$ km). These results for small planetesimal radii are dominated by these highly oblique impacts, and to verify the accuracy of these results, further work is needed in low-velocity, low-strength/strengthless, highly oblique impacts.
@@ 6 Discussion
We discuss our work in the context of the study by Nesvorný et al. (2021), who studied KBO binary formation with N-body soft sphere discrete element method (PKDGRAV). Nesvorný et al. (2021) found that the angular momentum of a pebble cloud prevents the rapid collapse into a single planetesimal. Instead, binary systems were more likely to form, with only a fraction of the pebble cloud being incorporated into the planetesimals. In this section we consider the fraction of pebble cloud angular momentum that must be removed to collapse a single planetesimal and place this value in context with the fraction of angular momentum that can be removed by escaping ejecta.
@@ 6.1 The Angular Momentum of Planetesimals Formed Via Streaming Instability
Equation 61 implies that the problem of angular momentum removal is worse in the outer solar system than in the inner solar system. In the outer solar system, due to the large fraction of angular momentum that must be lost, the angular momentum drain mechanism investigated in this paper would account for a large fraction of the required angular momentum drain, but would be insufficient to facilitate collapse into a single planetesimal without other mechanisms.
@@ 7 Summary
We find that the angular momentum drain mechanism proposed by Dobrovolskis and Burns (1984) operates on planetesimals that are embedded within a circumstellar disk and are experiencing impacts originating from a headwind. We characterize the efficiency of the process by computing the ratio of the angular momentum lost per unit projectile mass and the planetesimal’s spin angular momentum. We find that the angular momentum drain is efficient, with efficiency $\epsilon_{L}\sim 1$ for a wide range of parameters tested. These parameters include the projectile and planetesimal density being equal, a low-strength planetesimal with craters occurring in the strength regime, and headwind velocity being above $\sim 10$ m/s. If the projectile has greater density than the planetesimal, or if the impact occurs in the gravity regime, we find that the angular momentum drain increases, while a lower density projectile, decreased spin rate of the planetesimal, or very low headwind velocity result in decreased angular momentum drain. The expected total collisional mass is less than the mass of the planetesimal, as such we expect the total angular momentum drain to be 30 - 50% of the initial spin angular momentum of the planetesimal. With N-body simulations Nesvorný et al. (2021) found that the angular momentum of a pebble cloud prevents the collapse into a single planetesimal. Our study indicates that a pebble cloud with low initial angular momentum, with specific projectile population parameters, may collapse into single planetesimals.
We find the angular momentum transfer efficiency is most sensitive to headwind velocity and the ratio of planetesimal to projectile density. As planetesimal mass increases, the effects of gravitational focusing allow for greater angular momentum transfer from projectiles with a low headwind velocity due to accelerating the projectiles prior to impact. However, the mass loss ratio decreases as fewer particles reach the escape velocity. In the gravity regime, when projectile radius sets the $\pi_{2}$ parameter, we assume a projectile radius of $a_{pj}=1$ m, as the largest projectile radii and subsequently projectile mass will dominate the angular momentum transfer and mass loss. Assuming the projectiles originate from a disk filament that might be present due to streaming instability with a velocity gradient capable of generating the computed pebble cloud, we find no meaningful reduction of the angular momentum transfer efficiency with regard to spinning down an embedded planetesimal. Future studies could take into account the angular momentum that is inherent in a population of projectiles.
The impact related mechanism of angular momentum loss assists in removing a large portion of the excess angular momentum in a collapsing pebble cloud. However, this mechanism can not account for all of the angular momentum that needs to be removed, as such, additional sources of angular momentum drain may be necessary for the collapse of single-planetesimals. Pebble cloud collapse could lead preferentially to binary formation, as explored by Nesvorný et al. (2019, 2021). Alternatively, planetesimal formation itself could be inefficient with many pebble clouds failing to collapse, and only those that do collapse initially having low angular momentum.
@@
The python code used in this study is archived at https://arxiv.org/abs/2412.03533.
@@ Appendix A The Momentum and Angular Momentum Transfer
Assuming that the planetesimal is experiencing many collisions, if the distribution of projectiles has no angular momentum then the $m_{pj}{\bf r}\times{\bf u}_{pj}$ term in equation 69 will average to zero. We note that in equations 66 and 69, the relevant escaping ejecta momentum is computed in the frame rotating with the surface at the site of impact.
Up to this point, we have neglected gravitational focusing. Equations 62 and 63 describe the impact and ejecta curtain just prior to and after the impact. The projectile velocity ${\bf u}_{pj}$ refers to velocity of the projectile relative to the planetesimal at the moment of impact (in an inertial, not rotating frame) and is not equal to the velocity of the projectile at large distance from the planetesimal. As a consequence equation 68 describes the momentum transfer efficiency $\beta$ only for high velocity projectiles.